Woodlands Ltd., a paper and allied products manufacturer in Singapore, was seeking to gain a foothold in Indonesia. Toward that end, the company bought 40% of the outstanding ordinary shares of Jakar Timber Ltd. on January 1, 2016, for $400 million. It used equity method to account for the investment in Jakar. At the date of the purchase, the book value of Jakar’s net identifiable assets was $775 million. The fair value of Jakar’s inventory and equipment exceeded the book value by $50 million and $100 million, respectively. All other net identifiable assets had book value equal to fair value. All the inventory of Jakar at the beginning of 2016 was sold by year-end. Jakar used straight-line depreciation for its equipment which had 10 years remaining useful life and zero salvage value. Jakar reported net income of $140 million for the year ended December 31, 2016. Jakar paid a cash dividend of $30 million to all its shareholders. What amount should Woodlands report as income from its investment in Jakar for the year ended December 31, 2016?Woodlands Ltd., a paper and allied products manufacturer in Singapore, was seeking to gain a foothold in Indonesia. Toward that end, the company bought 40% of the outstanding ordinary shares of Jakar Timber Ltd. on January 1, 2016, for $400 million. It used equity method to account for the investment in Jakar. At the date of the purchase, the book value of Jakar’s net identifiable assets was $775 million. The fair value of Jakar’s inventory and equipment exceeded the book value by $50 million and $100 million, respectively. All other net identifiable assets had book value equal to fair value. All the inventory of Jakar at the beginning of 2016 was sold by year-end. Jakar used straight-line depreciation for its equipment which had 10 years remaining useful life and zero salvage value. Jakar reported net income of $140 million for the year ended December 31, 2016. Jakar paid a cash dividend of $30 million to all its shareholders. What amount should Woodlands report as income from its investment in Jakar for the year ended December 31, 2016?

A. $44 million.

B. $32 million.

C. $56 million.
Answer #1

Share of net income of Woodland from its investment in Jakar = 140 Million $ * 40% = 56 Milllion $

Adjustment of inventory = 40%*50 Million $*100% = 20 Million $

Adjustment for plant facilities = (100 Million $/10 years) * 40% = 4 Million $

Therefore share of net income in Jakar after adjustments = 56-20-4 = 32 Million $.

Therefore the correct answer is "B" : $32 million